OpenStudy (anonymous):

if you roll a dice, the chance of getting 6 is 1/6. If you roll the same dice 5 times, what is the probability of getting in all those 5 attempts only once a 6?

OpenStudy (zarkon):

use binomial dist

OpenStudy (anonymous):

so the answer is 59.81 percent thanks!

OpenStudy (zarkon):

if you want the probability of getting exactly one 6 in 5 rolls then the probability is \[{5\choose 1}\left(\frac{1}{6}\right)^1\left(\frac{5}{6}\right)^4=.4019\]